OpenStudy (samferguson):

Write the equation in vertex form for the parabola with vertex (0, 4) and focus (0, 6).

OpenStudy (3mar):

So at the beginning, it is a vertical parabola!  Can you create its equation?

OpenStudy (samferguson):

I am not sure how to create the equation at all

OpenStudy (3mar):

so Where are you stuck?

OpenStudy (samferguson):

I don't know where to start. Please give answer

OpenStudy (3mar):

It is not helpful for you to be given the answers without understanding how they are figured out! Step by step, I will give you through the answer!

OpenStudy (samferguson):

ok lets start

OpenStudy (3mar):

What is the general equation of the vertical parabola?

OpenStudy (samferguson):

\[y=ax ^{2}+b\]

OpenStudy (samferguson):

not sure if thats right

OpenStudy (3mar):

You mean that? \[\LARGE y=ax^2+bx+c\] or \[\LARGE y=a(x+b)^2+c\] which one are you familiar with?

OpenStudy (3mar):

@samferguson Do you follow?

OpenStudy (3mar):

Let me know where you are back!

OpenStudy (loser66):

When they give you the focus, then use it. (0, p) is a focus, so, \((x-h)^2=4p(y-k)\) when (h, k) is vertex.